OpenStudy (anonymous):

Write the slope-intercept form of the equation of (-6,7) and (4,-2) Show work

OpenStudy (anonymous):

Firstly you should find the slope.. Can you find that??

OpenStudy (anonymous):

the slope is -9/10

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

then would I do the y=m*x+b?

OpenStudy (anonymous):

Now use the following equation: \[y - y_1 = m(x - x_1)\] m is the slope and (x1, y1) is given as (-6, 7)..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

equation of a line of a given slope: y=mx+c where x is the x co-ordinate, y is the y co-ordinate and m is the slope.. here, you got the slope. the line passes through (-6,7) and (a,-2) choose any of them